# Estimating the Level of Dialectness Predicts Interannotator Agreement in Multi-dialect Arabic Datasets

### v1

@@ Abstract
On annotating multi-dialect Arabic datasets, it is common to randomly assign the samples across a pool of native Arabic speakers. Recent analyses recommended routing dialectal samples to native speakers of their respective dialects to build higher-quality datasets. However, automatically identifying the dialect of samples is hard. Moreover, the pool of annotators who are native speakers of specific Arabic dialects might be scarce. Arabic Level of Dialectness (ALDi) was recently introduced as a quantitative variable that measures how sentences diverge from Standard Arabic. On randomly assigning samples to annotators, we hypothesize that samples of higher ALDi scores are harder to label especially if they are written in dialects that the annotators do not speak. We test this by analyzing the relation between ALDi scores and the annotators’ agreement, on 15 public datasets having raw individual sample annotations for various sentence-classification tasks. We find strong evidence supporting our hypothesis for 11 of them. Consequently, we recommend prioritizing routing samples of high ALDi scores to native speakers of each sample’s dialect, for which the dialect could be automatically identified at higher accuracies.
@@ 1 Introduction
Despite this mutual unintelligibility, a common practice when developing datasets for multi-dialect Arabic NLP is to randomly recruit annotators without regard to their dialect. However, routing dialectal content to speakers of a different dialect for annotation or moderation can present real problems. For example, it has been shown to contribute to unjust online content moderation of DA Business for Social Responsibility (2022), and racially biased toxicity annotation in American English varieties Sap et al. (2022). Two recent studies of multi-dialect DA annotation showed that for annotating hate speech or sarcasm, respectively, annotators were more lenient (for hate speech) and more accurate (for sarcasm) when annotating sentences in their native dialect Bergman and Diab (2022); Abu Farha and Magdy (2022). The authors of both studies made the same recommendation for creating new Arabic datasets, namely to first identify the dialect of each sample and then route it to appropriate annotators.
@@ 2 Methodology
We study the impact of ALDi scores on the annotators’ agreement for publicly released Arabic datasets. We analyze datasets satisfying the following criteria:
@@
For each dataset, we compute the Arabic Level of Dialectness (ALDi) score for each annotated sample (sentence) using the Sentence-ALDi model Keleg et al. (2023), which returns a score from 0 (MSA/non-dialectal) to 1 (strongly dialectal). To investigate the effect of ALDi on annotator agreement, we bin the samples by their ALDi score into 10 bins of width 0.1. We compute % full agree, the percentage of samples in that bin for which all the annotators agreed on a single label. We employ Pearson’s correlation coefficient to analyze the relation between ALDi (represented by each bin’s midpoint ALDi score) and % full agree, and also report the slope of the best-fitting line as a measure of the effect size. [^4] The exact values of the slopes and correlation coefficients depend on the number of bins. However, we got similar qualitative results on using 4 or 20 equal-width bins. 10 bins are enough to check if trends are non-linear while keeping a reasonable number of samples in the smallest bins. We also fitted logistic regression (logreg) models using ALDi as a continuous variable and a binary outcome Full Agreement (Yes/No) for each sample. Both analysis tools reveal similar patterns (see Appendix §D) but the binning method provides useful additional visualization. As aforementioned, our initial hypothesis is that % full agree negatively correlates with high ALDi scores.
@@ 3 Results and Discussion
Inspecting the trends depicted in Figure 1, strong negative Pearson’s correlation coefficients exist for 8 out of the 12 datasets for the non-DI tasks (sentiment analysis, sarcasm, hate speech, and stance detection). Both the trends (quantified by the slope m) and the correlation coefficients for most of the tasks indicate that the percentage of samples for which all the annotators assign the same label decreases as the ALDi scores increase, often by a large margin. [^5] Refer to Appendix §F for a possible explanation of the unexpected trends of the ArSAS dataset. We notice different trends for DI that we elaborate on below.
@@ 4 Analysis of Trends by Class Label
A more nuanced analysis of the non-DI datasets can be done by splitting the samples according to their majority-vote labels (See Appendix §E).
@@
Conversely, the presence of specific words might be a strong cue for the label of the overall sentence, which might be useful for the positive classes such as (i.e.: Sarcastic, Positive/Negative, Obscene/Offensive). Consequently, annotators might be able to agree on a label for a sentence even if it contains some unintelligible segments. Consider the offensive sentence in Table 2. The presence of the MSA insult (you insect) is enough to guess that the sentence is offensive, even if the remaining segment is not fully intelligible. Similar high-ALDi samples are provided in Appendix §C. Lastly, note that agreeing on a label does not imply it is accurate in the case of relying on cues for annotation.
Morocco doesn’t know you, I don’t even know you, you insect
Table 2: An offensive example of a high ALDi score with a clear offensive MSA underlined segment.
@@ 5 Conclusion and Recommendation
Previous research recommended routing samples to native speakers of the different Arabic dialects for higher annotation quality. Our analysis indicates that a large proportion of 6 datasets are samples with ALDi scores < 0.1, which are expected to be MSA samples that can be routed to speakers of any Arabic dialect. Moreover, the lower agreement scores for samples with high ALDi scores show that extra care should be given to these samples. Dataset creators should prioritize routing them to native speakers of the dialects of these samples, which can be automatically identified at higher accuracy as these samples have more dialectal cues.
@@ Appendix B Code-mixing between MSA and DA
Appendix C Examples of High ALDi Samples
For samples with high ALDi scores, annotators might be looking for cues to make educated guesses, as elaborated in §4. Table C2 shows four examples, in which the annotators can use cues to label the whole sentence. It is hard to know whether the annotators relied on these cues to make their decision or not. We think the usage of cues is an intuitive explanation that needs further validation.
@@ Appendix C Examples of High ALDi Samples
Table C2: Four qualitative samples of high ALDi scores. The underlined segments represent the cues that the annotators might have used to choose a label even if they do not fully understand the sentence. Despite the presence of these cues, the annotators still disagreed on labeling the first two samples. We only provide translations for the underlined cue segments.
Appendix D Discussion about the Analysis
@@
Figure D1 demonstrates the probability of full agreement of each dataset, in addition to the coefficient of the ALDi score with its 95% confidence interval. For the 8 non-DI datasets with $Coef_{ALDi}<-0.2$, the coefficients can be considered to be statistically significant since the confidence interval does not include zero.
Both analysis tools achieve similar results. The same 8 non-DI datasets—ASAD, ArSarcasm-v1 (Sentiment/Sarcasm), Mawqif (Sentiment/Sarcasm/Stance), iSarcasm, and YTCB—have significantly strong negative correlation coefficients as in Figure 1, and statistically significant coefficients for the ALDi variable which are less than -0.2. However, binning the data allows for visualizing the % Full Agreement as a scatter plot, which can reveal whether the relation between ALDi and the agreement is linear or not, in addition to having a visual way for determining how well the best-fitting line models the data.
@@
Figure D1: For each dataset, plots show the estimated probability of Full-agreement according to each dataset’s fitted logistic regression model. Under each plot, the coefficient of ALDi with its 95% confidence interval is visualized. Nearly all datasets (marked with *) have confidence intervals that do not include zero, meaning the effect of ALDi is statistically significant at $p<0.05$. Negative coefficients indicate that higher ALDi scores predict lower agreement.
Appendix E Trends by Class Label
Figure E2: The trends for the classes of the Saracasm Detection datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure E3: The trends for the classes of the Sentiment Analysis datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure E4: The trends for the classes of the Offensive Text Classification and Hate Speech datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure E5: The trends for the Assertion and Expression labels of the ArSAS dataset, which represent 95% of the dataset samples. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure E6: The trends for the classes of Mawqif’s Stance dataset. Statistically significant correlation coefficients ($\rho$) are marked with *.
As mentioned in §4, Figures E2, E3, E4, E5, and E6 visualize the impact of ALDi on the annotator agreement after splitting the samples according to their majority-vote labels. We acknowledge that the number of samples in the bins for some classes is not enough to draw concrete conclusions (e.g.: samples with high ALDi scores for the Neutral class of the ArSAS, and Mawqif datasets as per Figure E3).
Appendix F The Rising Trend of ArSAS
The ArSAS dataset stands out as a dataset with a rising trend for the Speech Act Detection task and a falling trend for the Sentiment Analysis task. Samples of ArSAS were jointly annotated for their sentiment and speech act. Despite having 6 different speech acts, which would arguably make speech act detection harder than sentiment analysis, the Assertion and Expression classes represent 95% of the samples. Looking at their respective trends shown in Figure E5, the two acts show two different behaviors. Most of the assertive samples have ALDi scores <0.2 (arguably all are MSA ones). Moreover, the number of Assestion samples with high ALDi scores is not enough to estimate the % full agree for their respective bins. Conversely, the Expression act shows higher agreement as the ALDi score increases.

### v2

@@ Abstract
On annotating multi-dialect Arabic datasets, it is common to randomly assign the samples across a pool of native Arabic speakers. Recent analyses recommended routing dialectal samples to native speakers of their respective dialects to build higher-quality datasets. However, automatically identifying the dialect of samples is hard. Moreover, the pool of annotators who are native speakers of specific Arabic dialects might be scarce.
Arabic Level of Dialectness (ALDi) was recently introduced as a quantitative variable that measures how sentences diverge from Standard Arabic. On randomly assigning samples to annotators, we hypothesize that samples of higher ALDi scores are harder to label especially if they are written in dialects that the annotators do not speak. We test this by analyzing the relation between ALDi scores and the annotators’ agreement, on 15 public datasets having raw individual sample annotations for various sentence-classification tasks. We find strong evidence supporting our hypothesis for 11 of them. Consequently, we recommend prioritizing routing samples of high ALDi scores to native speakers of each sample’s dialect, for which the dialect could be automatically identified at higher accuracies.
@@ 1 Introduction
Despite this mutual unintelligibility, a common practice when developing datasets for multi-dialect Arabic NLP is to randomly recruit annotators without regard to their dialect.
However, routing dialectal content to speakers of a different dialect for annotation or moderation can present real problems. For example, it has been shown to contribute to unjust online content moderation of DA Business for Social Responsibility (2022), and racially biased toxicity annotation in American English varieties Sap et al. (2022). Two recent studies of multi-dialect DA annotation showed that for annotating hate speech or sarcasm, respectively, annotators were more lenient (for hate speech) and more accurate (for sarcasm) when annotating sentences in their native dialect Bergman and Diab (2022); Abu Farha and Magdy (2022).
The authors of both studies made the same recommendation for creating new Arabic datasets, namely to first identify the dialect of each sample and then route it to appropriate annotators.
@@ 2 Methodology
We study the impact of ALDi scores on the annotators’ agreement for publicly released Arabic datasets.
We analyze datasets satisfying the following criteria:
@@
For each dataset, we compute the Arabic Level of Dialectness (ALDi) score for each annotated sample (sentence) using the Sentence-ALDi model Keleg et al. (2023), which returns a score from 0 (MSA/non-dialectal) to 1 (strongly dialectal). To investigate the effect of ALDi on annotator agreement, we bin the samples by their ALDi score into 10 bins of width 0.1. We compute % full agree, the percentage of samples in that bin for which all the annotators agreed on a single label. We employ Pearson’s correlation coefficient to analyze the relation between ALDi (represented by each bin’s midpoint ALDi score) and % full agree, and also report the slope of the best-fitting line as a measure of the effect size [^4] The exact values of the slopes and correlation coefficients depend on the number of bins. However, we got similar qualitative results on using 4 or 20 equal-width bins. 10 bins are enough to check if trends are non-linear while keeping a reasonable number of samples in the smallest bins. We also fitted logistic regression (logreg) models using ALDi as a continuous variable and a binary outcome Full Agreement (Yes/No) for each sample. Both analysis tools reveal similar patterns (see Appendix §C) but the binning method provides useful additional visualization.. As aforementioned, our initial hypothesis is that % full agree negatively correlates with high ALDi scores.
@@ 3 Results and Discussion
YTCB Hatespeech Detection you insect 0.98 HateSpeech(3x)
you scabies 0.81 HateSpeech(3x)
Gulf’s colt 0.98 HateSpeech(3x)
# ??? # hahaha 0.94 Positive(2x) Negative(1x)
Table 2: Five qualitative samples of high ALDi scores. The underlined segments represent the cues that the annotators might have used to choose a label even if they do not fully understand the sentence. Despite the presence of these cues, the annotators still disagreed on labeling the last two samples. We only provide translations for the underlined cue segments.
@@
Inspecting the trends depicted in Figure 1, strong negative Pearson’s correlation coefficients exist for 8 out of the 12 datasets for the non-DI tasks (sentiment analysis, sarcasm, hate speech, and stance detection).
Both the trends (quantified by the slope m) and the correlation coefficients for most of the tasks indicate that the percentage of samples for which all the annotators assign the same label decreases as the ALDi scores increase, often by a large margin [^5] Refer to Appendix §E for a possible explanation of the unexpected trends of the ArSAS dataset.. We notice different trends for DI that we elaborate on below.
@@ 4 Analysis of Trends by Class Label
A more nuanced analysis of the non-DI datasets can be done by splitting the samples according to their majority-vote labels (See Appendix §D).
@@
Conversely, the presence of specific words might be a strong cue for the label of the overall sentence, which might be useful for the positive classes such as (i.e.: Sarcastic, Positive/Negative, Obscene/Offensive). Consequently, annotators might be able to agree on a label for a sentence even if it contains some unintelligible segments. Table B2 shows five examples, in which the annotators can use cues to label the whole sentence. Consider the first offensive sentence in Table LABEL:tab:qualitiative_seducated_guess_examples. The presence of the MSA insult (you insect) is enough to guess that the sentence is offensive, even if the remaining segment is not fully intelligible.
It is hard to know whether the annotators relied on these cues to make their decision or not. We think that the usage of cues is an intuitive explanation that needs further validation. We would also like to Lastly, note that agreeing on a label does not imply it is accurate, especially when in the case of relying on cues for annotation.
@@ 5 Conclusion and Recommendation
Previous research recommended routing samples to native speakers of the different Arabic dialects for higher annotation quality.
Our analysis indicates that a large proportion of 6 datasets are samples with ALDi scores < 0.1, which are expected to be MSA samples that can be routed to speakers of any Arabic dialect. Moreover, the lower agreement scores for samples with high ALDi scores show that extra care should be given to these samples. Dataset creators should prioritize routing them to native speakers of the dialects of these samples, which can be automatically identified at higher accuracy as these samples have more dialectal cues.
@@ Appendix B Code-mixing between MSA and DA
Table B2: Four qualitative samples of high ALDi scores. The underlined segments represent the cues that the annotators might have used to choose a label even if they do not fully understand the sentence. Despite the presence of these cues, the annotators still disagreed on labeling the first two samples. We only provide translations for the underlined cue segments.
Appendix C Discussion about the Analysis
@@
Figure C1 demonstrates the probability of full agreement of each dataset, in addition to the coefficient of the ALDi score with its 95% confidence interval. For the 8 non-DI datasets with $Coef_{ALDi}<-0.2$, the coefficients can be considered to be statistically significant since the confidence interval does not include zero.
Both analysis tools achieve similar results. The same 8 non-DI datasets—ASAD, ArSarcasm-v1 (Sentiment/Sarcasm), Mawqif (Sentiment/Sarcasm/Stance), iSarcasm, and YTCB—have significantly strong negative correlation coefficients as in Figure 1, and statistically significant coefficients for the ALDi variable which are less than -0.2. However,
binning the data allows for visualizing the % Full Agreement as a scatter plot, which can reveal whether the relation between ALDi and the agreement is linear or not, in addition to having a visual way for determining how well the best-fitting line models the data.
@@
Figure C1: For each dataset, plots show the estimated probability of Full-agreement according to each dataset’s fitted logistic regression model. Under each plot, the coefficient of ALDi with its 95% confidence interval is visualized. Nearly all datasets (marked with *) have confidence intervals that do not include zero, meaning the effect of ALDi is statistically significant at $p<0.05$. Negative coefficients indicate that higher ALDi scores predict lower agreement.
Appendix D Trends by Class Label
Figure D2: The trends for the classes of the Saracasm Detection datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure D3: The trends for the classes of the Sentiment Analysis datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure D4: The trends for the classes of the Offensive Text Classification and Hate Speech datasets. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure D5: The trends for the Assertion and Expression labels of the ArSAS dataset, which represent 95% of the dataset samples. Statistically significant correlation coefficients ($\rho$) are marked with *.
Figure D6: The trends for the classes of Mawqif’s Stance dataset. Statistically significant correlation coefficients ($\rho$) are marked with *.
As mentioned in §4, Figures D2, D3, D4, D5, and D6 visualize the impact of ALDi on the annotator agreement after splitting the samples according to their majority-vote labels. We acknowledge that the number of samples in the bins for some classes is not enough to draw concrete conclusions (e.g.: samples with high ALDi scores for the Neutral class of the ArSAS, and Mawqif datasets as per Figure D3).
Appendix E The Rising Trend of ArSAS
The ArSAS dataset stands out as a dataset with a rising trend for the Speech Act Detection task and a falling trend for the Sentiment Analysis task. Samples of ArSAS were jointly annotated for their sentiment and speech act. Despite having 6 different speech acts, which would arguably make speech act detection harder than sentiment analysis, the Assertion and Expression classes represent 95% of the samples. Looking at their respective trends shown in Figure D5, the two acts show two different behaviors. Most of the assertive samples have ALDi scores <0.2 (arguably all are MSA ones). Moreover, the number of Assestion samples with high ALDi scores is not enough to estimate the % full agree for their respective bins. Conversely, the Expression act shows higher agreement as the ALDi score increases.